OpenStudy (anonymous):

Please explain? And don't give me the answer. I just need some help. Solve: 2(x + 8) − (x + 1) = 20

OpenStudy (anonymous):

Lol....Use Distributive LAw to Open the Brackets and Solve

OpenStudy (anonymous):

2x + 16 - x - 1 = 20 nw Solve for x

OpenStudy (anonymous):

C(x+y)=Cx+Cy. Once you get rid of the first parentheses, you have to distribute the negative sign as a -1 to get rid of the second parentheses and then combine like terms and isolate x.

OpenStudy (anonymous):